OpenStudy (anonymous):

23 tafara27 Compare taking a root and rising to a power. Support your observations with your own examples.

OpenStudy (unklerhaukus):

\[23\]\[\sqrt{23}\approx 4.8\]\[23^2=529\]

Parth (parthkohli):

Hmm, I believe that he had copied his SmartScore.

Parth (parthkohli):

he copied his smartscore and username.

OpenStudy (unklerhaukus):

\[\sqrt[3]{23}\approx2.8\] \[23^3=12167\]

OpenStudy (anonymous):

thanks @UnkleRhaukus so how you compare it ?

OpenStudy (unklerhaukus):

so we have found when taking an nth root, - the number is small and taking the nth power , the number is bigger where n>1

OpenStudy (unklerhaukus):

smaller*

OpenStudy (anonymous):

hey @UnkleRhaukus I have a ????What method of solving quadratic equations did you learn in Algebra ? What method do you prefer, and in what cases?

OpenStudy (unklerhaukus):

\[ax^2+bx+c=0\]if i can factor a quadratic equations in my head i will, otherwise i'd use the quadratic formula \[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]